OpenStudy (anonymous):

Determine the function f satisfying the given conditions. f ' (x) = e^x/3 f (0) = 9 f (x) = A e^Bx + C

OpenStudy (loser66):

@SithsAndGiggles for sure you get the correct answer, hihihi

OpenStudy (anonymous):

\[\begin{align*} f'(x)&=e^{x/3}\\ \int f'(x)~dx&=\int e^{x/3}~dx\\ f(x)&=\cdots \end{align*}\] Plug in the initial condition and solve for the constant of integration, etc.

OpenStudy (anonymous):

A=3, B=1/3, C=6?